• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name tony

    Log entry time 08:05:37 on May 7,2000

    Entry number 40667

    keyword=Owl Shift Summary

    Shift Summary
    00:30 Beginning shift without beam, phoned MCC for update:
    Problem in Bldg 1 of the N. linac

    01:25 CW returs to the hall (max current ~53 uA)
    Run# 2089 Started 01:37

    05:05 MCC down for 5min QE measurement
    Ended run # 2091 (72k)

    05:45 RF Trip in S. linac ; ended run# 2092 (75k)

    06:45 Major S. linac crash ; down for a while
    ended run# 2093 (124k)

    07:25 CW restored to hall; more production running